# FSPGD: Rethinking Black-box Attacks on Semantic Segmentation

### v1

@@ Abstract
Transferability, the ability of adversarial examples crafted for one model to deceive other models, is crucial for black-box attacks. Despite advancements in attack methods for semantic segmentation, transferability remains limited, reducing their effectiveness in real-world applications. To address this, we introduce the Feature Similarity Projected Gradient Descent (FSPGD) attack, a novel black-box approach that enhances both attack performance and transferability. Unlike conventional segmentation attacks that rely on output predictions for gradient calculation, FSPGD computes gradients from intermediate layer features. Specifically, our method introduces a loss function that targets local information by comparing features between clean images and adversarial examples, while also disrupting contextual information by accounting for spatial relationships between objects. Experiments on Pascal VOC 2012 and Cityscapes datasets demonstrate that FSPGD achieves superior transferability and attack performance, establishing a new state-of-the-art benchmark. Code is available at https://github.com/KU-AIVS/FSPGD.
@@ 1 Introduction
To overcome this problem, various adversarial attack methods [16, 25, 51, 1, 5, 4, 26] specifically designed for semantic segmentation have been introduced. While these methods show fine attack performance in semantic segmentation, they have not yet fully overcome the challenges of transferability. In this study, we analyze the reasons for the weak transferability of existing methods and identify the following causes: conventional methods usually calculate gradients and generate perturbations by using the output predictions of the source model. This approaches exhibit strong attack performance only on the source model but fail to achieve similar performance on new target models. This limitation arises because these methods only consider pixel-wise predictions and do not effectively attack contextual information, i.e. the spatial relationships between objects, which is a critical factor in semantic segmentation.
To address this problem, this paper proposes a novel black-box attack method, called the Feature Similarity Projected Gradient Descent (FSPGD) attack, which demonstrates strong attack performance and significant transferability. Unlike existing segmentation attack methods that rely solely on output predictions from the source model to compute gradients, the proposed method calculates gradients by leveraging features extracted from the intermediate layer. Specifically, we develop a novel loss function that targets local information by comparing features between clean images and adversarial examples, while also disrupting contextual information by leveraging spatial relationships between objects within the image. To validate the superiority of the proposed method, we present comprehensive experimental results across a variety of models, such as PSPNet-ResNet50 [55], PSPNet-ResNet101 [55], DeepLabv3-ResNet50 [6], DeepLabv3-ResNet101 [6], and FCN-VGG16 [32], trained by using Pascal VOC 2012[11] and Cityscapes [8] datasets. Moreover, a series of ablation studies are conducted to highlight the robust generalization capabilities of the proposed method. Quantitative evaluations clearly show that the proposed method not only achieves strong attack performance but also surpasses conventional methods in transferability, setting a new state-of-the-art benchmark. Our contribution can be summarized as follows:
@@
Figure 1: Visualization of the feature similarity. We show a feature similarity map using the features of the bicycle wheel area (red box) as the reference feature. In conventional methods, high feature similarity is observed with other bicycle wheels (yellow and blue boxes), whereas in the proposed method, feature similarity is notably reduced. (a) Clean image, (b) PGD [34] (c) SegPGD [16], (d) CosPGD [1], and (e) FSPGD (Ours).
@@ 3.1 Motivation
To validate our hypothesis, we conducted experiments to visualize feature similarity in the intermediate layer, as illustrated in Fig. 1. Using the feature vector of the bicycle wheel region (red box) as a reference, we generated a similarity map across other areas. Adversarial images are made by using DeepLabV3-ResNet50 as the source model and DeepLabV3-ResNet101 as the target model. In Fig. 1(a), the clean image shows that the reference feature aligns with those of other bicycle wheel regions (yellow and blue boxes), indicating that the network generates similar features for objects of the same class, even when spatially separated. Despite the attack, as described in Figs. 1(b), (c), and (d), conventional methods still show similar features in the target model. Objects of the same class continue to show similar features, leading to weak attack performance with predictions nearly identical to those of the clean image, demonstrating low transferability in conventional methods. In contrast, the proposed method considers spatial relationships, leading to feature dissimilarity between wheel regions (red, yellow, and blue boxes). Additional examples of similarity maps are provided in the supplementary material.
@@ 3.2 Methodology
By combining Eqs. 4 and 8, we make our objective function L as follows:
$L=\lambda_{t}L_{ex}+(1-\lambda_{t})L_{in},$ (9)
@@
Table 1: Attack performance comparison on Pascal VOC 2012 and Cityscapes in terms of mIoU. We set networks containing Res50 encoder as the source model. Lower mIoU means better performance and bold numbers denote the best mIoU values for each experimental setup.
@@
Target Models Source PSP DV3 FCN Source PSP DV3 FCN
Model Res101 Res101 VGG16 model Res101 Res101 VGG16
Source Clean Images 80.22 / 80.18 78.39 82.88 59.80 64.62 / 65.90 65.65 67.16 57.00
PSP PGD [34] 7.72 54.73 59.41 45.70 1.70 21.43 22.00 27.37
Res50 SegPGD [16] 5.41 54.10 58.95 45.43 1.13 21.56 23.02 28.02
CosPGD [1] 1.84 56.63 64.37 45.99 0.09 25.47 28.14 28.38
FSPGD (Ours) 3.39 22.24 16.84 19.81 0.93 4.98 3.21 10.92
DV3 PGD [34] 9.74 52.96 56.35 46.39 1.86 25.11 25.21 28.19
Res50 SegPGD [16] 7.26 52.05 56.50 46.23 1.28 25.54 25.36 28.57
CosPGD [1] 1.67 56.82 61.36 45.94 0.05 26.89 28.86 27.89
FSPGD (Ours) 3.44 21.89 16.57 19.36 1.14 6.04 3.79 10.89
Table 2: Attack performance comparison on Pascal VOC 2012 and Cityscapes in terms of mIoU. We set networks containing Res101 encoder as the source model. Lower mIoU means better performance and bold numbers denote the best mIoU values for each experimental setup.
[…]
Target Models Source PSP DV3 FCN Source PSP DV3 FCN
Model Res50 Res50 VGG16 model Res50 Res50 VGG16
Source Clean Images 78.39 / 82.88 80.22 80.18 59.80 65.65 / 67.16 64.62 65.90 57.00
PSP PGD [34] 10.13 55.93 55.39 47.25 1.89 12.74 14.56 26.43
Res101 SegPGD [16] 7.31 53.56 54.03 46.26 0.83 13.00 14.23 26.53
CosPGD [1] 2.87 57.54 58.50 47.05 0.06 16.32 17.37 27.10
FSPGD (Ours) 2.99 12.57 13.65 21.31 2.16 2.41 3.20 11.88
DV3 PGD [34] 9.75 56.36 55.54 47.48 2.18 18.37 18.10 27.56
Res101 SegPGD [16] 7.18 54.47 53.96 46.53 1.18 18.34 17.82 27.83
CosPGD [1] 2.73 58.83 58.84 47.25 0.02 20.84 20.69 27.59
FSPGD (Ours) 3.28 11.44 13.47 21.67 2.32 2.34 2.92 11.65
4 Experiment
4.1 Experimental Setup
Datasets. We use two popular semantic segmentation datasets in our experiments: PASCAL VOC 2012 [11] and Cityscapes [8]. The VOC dataset includes 20 object classes and one background class, containing 1,464 images for training and 1,499 for validation. Following the standard protocol [19], the training set is expanded to 10,582 images. The Cityscapes dataset, focused on urban scene understanding, comprises 19 categories with high-quality pixel-level annotations, including 2,975 images for training and 500 for validation. In our experiments, attack performance is evaluated using the validation set of each dataset.
Models. In this paper, we employ popular semantic segmentation models, i.e. PSPNet-ResNet50 [55], PSPNet-ResNet101 [55], DeepLabv3-ResNet50 [6], DeepLabv3-ResNet101 [6], and FCN-VGG16 [32], for our source and target models. We conduct cross-validation by alternating source and target models to demonstrate the transferability of the proposed method. For instance, when DeepLabv3-ResNet50 is used as the source model, we measure attack performance on DeepLabv3-ResNet101, PSPNet-ResNet101, and FCN-VGG16.
Parameters. Each comparison experiment follows the $l_{\infty}$-norm, setting the maximum perturbation value $\epsilon$ to 8/255 The step size $\alpha$ is set to 2/255. The proposed method has a user parameter $\tau$ which acts the threshold value in Eq. 7. In our experiments, we set $\tau$ value as cos($\pi/3$). The reason we set the threshold value as a cosine value is as follows: since $\textrm{M}(p,q)$ is calculated through the inner product of two vectors with a magnitude of 1, its value represents the cosine of the angle $\theta$ between two vectors. Therefore, we choose the threshold value based on the cosine value.
4.2 Experimental Results
We first compare the attack performance on conventional methods [1, 16, 34] with the proposed method. The experimental results are summarized in Tables 1 and 2. PSPResX and DV3ResX indicate the PSPNet [55] and DeepLabV3 [6] with ResNet50 [20] (or ResNet101 [20]) encoder, respectively. The proposed method shows high attack performance on the source model compared to conventional methods, excluding CosPGD [1] which is designed for white-box attack. To evaluate transferability, we measure mIoU on various target models. In this study, we select target models such that the encoders (e.g. ResNet50 and ResNet101) do not overlap between source and target models. As shown in Tables 1 and 2, the proposed method exhibits significantly superior transferability compared to conventional methods. In particular, it shows strong attack performance not only on target models using ResNet-based encoders but also on substantially different models, such as FCN with VGG16. These results indicate that the proposed method is better suited for real-world scenarios compared to traditional methods. Due to page limitations, we compare the performance of only a few conventional methods in Tables 1 and 2. Additional experimental results comparing a wider range of conventional methods are described in the supplementary material.
Figure 3: Visualization of clean image, attacked images, and output predictions. Deeplabv3-Res50 is used as the source model and Deeplabv3-Res101 (second row) and PSPNet-Res101 (third row) are used as target models. (a) Clean image, (b) PGD [34], (c) SegPGD [16], (d) CosPGD [1], (e) FSPGD (Ours).
For qualitative evaluation, we visualize adversarial examples along with their corresponding prediction results. In our experiments, we set DeepLabV3-ResNet50 as the source model and DeepLabV3-ResNet101 as the target model. As shown in Fig. 3, prediction results of conventional methods are similar to the results on clean images, indicating weak transferability. In contrast, the proposed method successfully attacks target models, demonstrating strong transferability. Based on these results, we conclude that the proposed method achieves the state-of-the-art transferability performance. Additional images of attack results are provided in the supplementary material.
4.3 Ablation Studies
The proposed method incorporates a user-defined variable $\tau$ for binarizing M. To determine the optimal $\tau$ value, we conduct ablation studies on Pascal VOC 2012 dataset. Fig. 4 presents a graph of mIoU performance based on different $\tau$ values. To select the $\tau$ value that maximizes transferability, we conduct experiments by setting the source and target models differently. As shown in Fig. 4, the average mIoU value is the lowest when $\tau$ is set to cos($\pi$/3), leading us to select cos($\pi$/3) in our study. It is noteworthy that, as shown in Tables 1,  2, and Fig. 4, the proposed method still shows superior performance compared to conventional methods, regardless of the $\tau$ value. Therefore, we believe that, despite having a user-defined parameter, the proposed method offers the advantage of superior performance compared to existing methods.
Meanwhile, the proposed loss function consists of two loss terms, $L_{ex}$ and $L_{in}$. To examine the effect of each loss term, we conduct an ablation study on the following three combinations: i) using $L_{ex}$ loss only, ii) using $L_{in}$ loss only, and iii) using $L_{ex}$ + $L_{ex}$ without $\lambda_{t}$ value. First, as shown in Fig. 5, using only $L_{ex}$ shows strong attack performance on the source model and achieves a fine transferability performance. In contrast, using only $L_{in}$ results in poor attack performance. This is because $L_{in}$ performs the attack by reducing internal similarity within $f_{a}$ without comparison to $f_{x}$, making it less effective when used alone. On the other hand, when $L_{ex}$ and $L_{in}$ are used together without employing $\lambda_{t}$ value, it shows similar or slightly weak performance than when using $L_{ex}$ loss alone. We believe that the effectiveness of each loss term is compromised when $L_{ex}$ and $L_{in}$ are simultaneously applied in equal proportions. However, when using the $\lambda_{t}$ value, the attack performance consistently improves compared to using only $L_{ex}$ (case i) or simply adding both $L_{ex}$ and $L_{in}$ (case iii). Therefore, in this study, we perform the attack by adjusting the balance of the two losses at each iteration using $\lambda_{t}$.
Furthermore, we conduct ablation studies to identify the most effective feature extraction locations within the source model. The results, illustrated in Fig. 6, display mIoU performance across different extraction points. The layer names specify feature map extraction locations within the encoder (specifically, ResNet50 and ResNet101). For example, the notation ‘layer name 2_1’ refers to the output of the first block within the ResNet Conv2_1 layer. As shown in the graph, selecting an intermediate layer within the encoder for the attack yields better results compared to using either the early layers or the later layers. Based on these findings, we identify the extraction points with the lowest average mIoU for our experiments, which are layer name 3_2 in ResNet50 and layer name 3_10 in ResNet101.
Figure 4: mIoU performance across different $\tau$ values. (S) and (T) indicate the source and target models, respectively.
Figure 6: mIoU performance across different locations of intermediate layers. (S) and (T) indicate the source and target models, respectively.
5 Limitations
6 Conclusion
In this paper, we identify key limitations in existing segmentation attack methods and conduct an in-depth analysis of the underlying causes. Based on these observations, we develop and introduce a novel segmentation attack method, called Feature Similarity Projected Gradient Descent (FSPGD), specifically designed to enhance both attack performance and transferability. The proposed FSPGD method demonstrates notable improvements over conventional methods, not only in terms of attack efficacy but also in transferability across different model architectures. To validate the superiority of FSPGD, we conduct extensive experiments across various source and target models, including PSPNet-ResNet50, PSPNet-ResNet101, DeepLabv3-ResNet50, DeepLabv3-ResNet101, and FCN-VGG16. Future work will aim to further optimize the parameter settings of FSPGD to enhance its robustness and adaptability across various model configurations. Additionally, we plan to explore a more automated approach for parameter optimization, which would allow the method to achieve optimal results efficiently across a diverse set of models, thus broadening its applicability in real-world scenarios.
[…]
Supplementary Material
A Visualization of Feature Similarity
To further validate the motivation described in Sec. 3.1 , we performed visualizations on a broader variety of images. Figs. 1 and  2 present experimental results on the Pascal VOC 2012 dataset, while Figs. 3 and  4 show results on the Cityscapes dataset. As seen in the figures, conventional methods maintain the similarity of features within the same class even after performing an attack, leading to poor attack performance on new target models. In contrast, the proposed method reduces feature similarity and exhibits superior attack performance compared to conventional methods.
B Extended Experimental Results
To further prove the superiority of the proposed method, we conducted comparative experiments with various conventional methods [34, 16, 1, 51, 31, 52, 10]. To evaluate the transferability of the attack methods, we designed the experiments with non-overlapping encoders for the source model and target model. As shown in Tables 1 and 2, the proposed method achieves the best performance among black-box attack methods on the source model (CosPGD [1] is a white-box attack method) and shows superior attack performance on target models compared to existing methods.
Table 1: Attack performance comparison on Pascal VOC 2012 in terms of mIoU. We set networks containing ResNet50 encoder as the source model. Lower mIoU means better performance and bold numbers denote the best mIoU values for each experimental setup.
  Target Models Source Model PSPRes101 DVRes101 FCNVGG16
Source Models Clean Images 80.22/80.18 78.39 82.88 59.80
@@ B Extended Experimental Results
Table 2: Attack performance comparison on Pascal VOC 2012 in terms of mIoU. We set networks containing ResNet101 encoder as the source model. Lower mIoU means better performance and bold numbers denote the best mIoU values for each experimental setup.
  Target Models Source Model PSPRes50 DVRes50 FCNVGG16
Source Models Clean Images 78.39/82.88 80.22 80.18 59.80

### v2

@@ Abstract
Transferability, the ability of adversarial examples crafted for one model to deceive other models, is crucial for black-box attacks. Despite advancements in attack methods for semantic segmentation, transferability remains limited, reducing their effectiveness in real-world applications. To address this, we introduce the Feature Similarity Projected Gradient Descent (FSPGD) attack, a novel black-box approach that enhances both attack performance and transferability. Unlike conventional segmentation attacks that rely on output predictions for gradient calculation, FSPGD computes gradients from intermediate layer features. Specifically, our method introduces a loss function that targets local information by comparing features between clean images and adversarial examples, while also disrupting contextual information by accounting for spatial relationships between objects. Experiments on Pascal VOC 2012 and Cityscapes datasets demonstrate that FSPGD achieves superior transferability and attack performance, establishing a new state-of-the-art benchmark. Code is available at https://anonymous.4open.science/r/FSPGD/README.md.
@@ 1 Introduction
To overcome this problem, various adversarial attack methods [17, 26, 53, 1, 5, 4, 27] specifically designed for semantic segmentation have been introduced. While these methods show fine attack performance in semantic segmentation, they have not yet fully overcome the challenges of transferability. In this study, we analyze the reasons for the weak transferability of existing methods and identify the following causes: conventional methods usually calculate gradients and generate perturbations by using the output predictions of the source model. This approach exhibits strong attack performance only on the source model but fail to achieve similar performance on new target models. This limitation arises because these methods only consider pixel-wise predictions and do not effectively attack contextual information, i.e. the spatial relationships between objects, which is a critical factor in semantic segmentation.
To address this problem, this paper proposes a novel black-box attack method, called the Feature Similarity Projected Gradient Descent (FSPGD) attack, which demonstrates strong attack performance and significant transferability. Unlike existing segmentation attack methods that rely solely on output predictions from the source model to compute gradients, the proposed method calculates gradients by leveraging features extracted from the intermediate layer. Specifically, we develop a novel loss function that targets local information by comparing features between clean images and adversarial examples, while also disrupting contextual information by leveraging spatial relationships between objects within the image. To validate the superiority of the proposed method, we present extensive experimental results across a variety of models, such as PSPNet-ResNet50 [58], DeepLabv3-ResNet50 [6], SegFormer-MiT B0 [55], and Mask2Former-Swin S [8]. Moreover, a series of ablation studies are conducted to highlight the robust generalization capabilities of the proposed method. Quantitative evaluations clearly show that the proposed method not only achieves strong attack performance but also surpasses conventional methods in transferability, setting a new state-of-the-art benchmark. Our contribution can be summarized as follows:
@@
Figure 1: Visualization of the feature similarity. We show a feature similarity map using the features of the bicycle wheels area (red box) as the reference feature. In conventional methods, high feature similarity is observed with other bicycle wheels (yellow and blue boxes), whereas in the proposed method, feature similarity is notably reduced. (a) Clean image, (b) PGD [36] (c) SegPGD [17], (d) CosPGD [1], and (e) FSPGD (Ours).
@@ 3.1 Motivation
To validate our hypothesis, we conducted experiments to visualize feature similarity in the intermediate layer, as depicted in Fig. 1. Using the feature vector of the bicycle wheel region (red box) as a reference feature, we generated a map comparing feature similarity with other areas, using DeepLabV3-ResNet50 as the source model and DeepLabV3-ResNet101 as the target model. As shown in Fig. 1(a), the clean image reveals that the reference feature is similar to those of other bicycle wheel regions (yellow and blue boxes), indicating that the network generates similar features for objects with the same class, even when they are spatially separated. Despite the attack, as shown in Figs. 1(b), (c), and (d), conventional methods still produce similar features in the target model. In other words, objects with the same class continue to exhibit similar features, leading to weak attack performance (producing predictions nearly identical to those for the clean image); these results show the low transferability in conventional methods. In contrast, the proposed method performs the attack by accounting for spatial relationships, resulting in feature dissimilarity between wheel regions (red, yellow, and blue boxes). Consequently, the proposed method achieves better attack performance and demonstrates superior transferability compared to conventional methods.
@@ 3.2 Methodology
Table 1: Attack performance comparison on Pascal VOC 2012 in terms of mIoU. Lower mIoU means better performance and bold numbers denote the best mIoU values for each experimental setup
@@
Source Models Attack Method Source Model PSPRes101 DV3Res101 FCNVGG16
Clean Images 80.22/80.18 78.39 82.88 59.80
@@
Table 2: Attack performance comparison on Cityscapes in terms of mIoU. Lower mIoU means better performance and bold numbers denote the best mIoU values for each experimental setup
[…]
Source Models Attack Source PSP DV3 PSP DV3 Mask2Former
Method Model Res50 Res50 Res101 Res101 Swin-S
Clean Images 60.58 64.62 65.65 65.90 67.16 68.24
SegFormer PGD [36] 1.06 29.94 36.07 31.99 38.25 48.43
MiT-B0 SegPGD [17] 0.38 28.45 34.56 29.28 36.38 49.54
CosPGD [1] 0.00 29.98 35.92 32.19 37.72 51.51
DAG [53] 50.92 20.84 33.73 32.71 28.77 55.21
NI [32] 2.06 30.27 37.63 30.95 38.24 43.75
DI [54] 9.13 41.92 45.85 43.10 48.06 46.78
TI [11] 7.66 50.60 52.77 52.25 55.88 55.59
FSPGD (Ours) 1.33 10.09 14.57 21.16 22.06 39.92
Source Models Attack Source PSP DV3 PSP DV3 SegFormer
Method Model Res50 Res50 Res101 Res101 MiT-B0
Clean Images 68.24 64.62 65.65 65.90 67.16 60.58
Mask2Former PGD [36] 0.45 39.41 45.25 42.15 48.35 49.30
Swin-S SegPGD [17] 0.30 39.97 45.07 42.29 48.96 49.40
CosPGD [1] 0.17 39.56 45.23 42.36 47.43 49.37
DAG [53] 65.59 30.69 42.06 32.76 39.42 54.23
NI [32] 0.17 42.76 49.41 45.06 50.00 45.87
DI [54] 3.53 50.34 53.67 53.16 56.59 50.85
TI [11] 0.85 56.81 59.74 59.95 62.69 59.74
FSPGD (Ours) 2.20 15.57 18.00 24.29 25.96 36.87
By combining Eqs. 4 and 8, we make our objective function L as follows:
$L=\lambda_{t}L_{ex}+(1-\lambda_{t})L_{in},$ (9)
4 Experiment
Figure 3: Visualization of experimental results. DV3Res50 is used as the source model and images of first column are clean images and adversarial examples generated by PGD [36], SegPGD [17], CosPGD [1], and FSPGD (Ours). second column is ground truth of input images. And other columns are predictions of target models.
4.1 Experimental Setup
Datasets. We use two popular semantic segmentation datasets in our experiments: PASCAL VOC 2012 [12], Cityscapes [9]. The VOC dataset includes 20 object classes and one background class, containing 1,464 images for training and 1,499 for validation. Following the standard protocol [20], the training set is expanded to 10,582 images. The Cityscapes dataset, focused on urban scene understanding, comprises 19 categories with high-quality pixel-level annotations, including 2,975 images for training and 500 for validation. In our experiments, attack performance is evaluated using the validation set of each dataset.
Models. In this paper, we employ popular semantic segmentation models, i.e. PSPNet-ResNet50, DeepLabv3-ResNet50 [6], SegFormer-MiT B0 [55], and Mask2Former-Swin S [8] as our source and target models, with FCN-VGG16 [34] additionally used as target model. We conduct cross-validation by alternating source and target models to demonstrate the transferability of the proposed method. For instance, when PSPNet-ResNet50 is used as the source model, we measure attack performance on DeepLabv3-ResNet101, PSPNet-ResNet101, FCN-VGG16.
Parameters. Each comparison experiment follows the $l_{\infty}$-norm, setting the maximum perturbation value $\epsilon$ to 8/255. The step size $\alpha$ is set to 2/255 and the total iteration $T$ is set to 20. The proposed method has a user parameter $\tau$ which acts the threshold value in Eq. 7. In our experiments, we set $\tau$ value as cos($\pi/3$). The reason we set the threshold value as a cosine value is as follows: since $\textrm{M}(p,q)$ is calculated through the inner product of two vectors with a magnitude of 1, its value represents the cosine of the angle $\theta$ between two vectors. Therefore, we choose the threshold value based on the cosine value.
4.2 Experimental Results
We first compare the attack performance on conventional methods [1, 17, 36, 53, 32, 54, 11] with the proposed method. The experimental results are summarized in Tables 1 and  2. PSPResX and DV3ResX indicate the PSPNet [58] and DeepLabV3 [6] with ResNet50 [21] (or ResNet101 [21]) encoder, respectively. In the Cityscapes, SegFormer [55] with a MiT-B0 [55] encoder and Mask2Former [8] with a Swin-S [33] encoder are used as transformer-based source models. The proposed method shows high attack performance on the source model compared to conventional methods, excluding CosPGD [1] which is designed for white-box attack. To evaluate transferability, we measure mIoU on various target models. In this study, we select target models such that the encoders (e.g. ResNet50 and ResNet101) do not overlap between source and target models. As shown in Tables 1 and 2, the proposed method exhibits significantly superior transferability compared to conventional methods. In particular, it shows strong attack performance not only on target models using ResNet-based encoders but also on substantially different models based on transformer in Table 2. These results indicate that the proposed method is better suited for real-world scenarios compared to traditional methods. Due to page limitations, we compare the performance of only a few source models in Tables 1 and  2. Additional experimental results comparing a wider range of conventional methods are described in the supplementary material.
For qualitative evaluation, we visualize adversarial examples along with their corresponding prediction results. In our experiments, we set DeepLabV3-ResNet50 as the source model. PSPNet with Resnet50 (and Resnet101)and DeepLabV3 with Resnet50 (and Resent101) set as the target model. As shown in Fig. 3, prediction results of conventional methods are similar to the results on clean images, indicating weak transferability. In contrast, the proposed method successfully attacks target models, demonstrating strong transferability. Based on these results, we conclude that the proposed method achieves the state-of-the-art transferability performance. Additional images of attack results are provided in the supplementary material.
4.3 Ablation Studies
The proposed method incorporates a user-defined variable $\tau$ for binarizing M. To determine the optimal $\tau$ value, we conduct ablation studies on Pascal VOC 2012 dataset. To select the $\tau$ value that maximizes transferability, we conduct experiments with all other variables fixed by setting $\tau$ to $cos(\pi/3)$, $cos(\pi/4)$, and $cos(\pi/6)$ and Fig. 4 presents the results. Since the average mIoU value was the lowest when $\tau$ was set to $cos(\pi/3)$, we selected $cos(\pi/3)$ in our study. As shown in Table 1 and Fig. 4, the proposed method outperforms existing methods, regardless of the $\tau$ value. Therefore, we believe that, despite having a user-defined parameter, the proposed method offers the advantage of superior performance compared to existing methods. The proposed loss function consists of two components: $L_{ex}$ and $L_{in}$. To evaluate the effect of each loss term, we conduct an ablation study with five different configurations. First, we consider using only the external loss term $L_{ex}$, and second, using only the internal loss term $L_{in}$. Third, we explore a fixed weight combination of the two losses, where $L_{ex}$ + $\lambda_{t}L_{in}$ is used with a constant $\lambda$ (e.g. 0.1, 0.5, 1.0). Finally, we employ an adaptive weighting strategy in which both losses are dynamically adjusted using $\lambda$ at each iteration, following the formulation $\lambda_{t}L_{ex}$ + $(\lambda_{t}-1)L_{in}$. As shown in Fig. 5, we observe that the experiment using the dynamic lambda strategy achieved better attack performance compared to other loss combinations. Hence, we employ the dynamic lambda strategy.
Furthermore, we conduct ablation studies to determine the most effective feature extraction layers in various source models. Trough conducting extensive experiments, we determine the optimal layers are Layer 2 of Conv3_x in ResNet-50, Layer 1 of Transformer block 1 in MiT-B0, and Layer 1 of Stage 2 in Swin-S. Detailed results are provided in the Supplemental Material.
5 Limitations
6 Conclusion
In this paper, we identify key limitations in existing segmentation attack methods and conduct an in-depth analysis of the underlying causes. Based on these observations, we develop and introduce a novel segmentation attack method, called Feature Similarity Projected Gradient Descent (FSPGD), specifically designed to enhance both attack performance and transferability. The proposed FSPGD method demonstrates notable improvements over conventional methods, not only in terms of attack efficacy but also in transferability across different model architectures. Future work will aim to further optimize the parameter settings of FSPGD to enhance its robustness and adaptability across various model configurations. Additionally, we plan to explore a more automated approach for parameter optimization, which would allow the method to achieve optimal results efficiently across a diverse set of models, thus broadening its applicability in real-world scenarios.
[…]
Supplementary Material
A Visualization of Feature Similarity
To further validate the motivation described in Sec. 3.1 , we performed visualizations on a broader variety of images. Figs. 1 and  2 present experimental results on the Pascal VOC 2012 dataset, while Figs. 3 and Figs. 4 show results on the Cityscapes dataset. As seen in the figures, conventional methods maintain the similarity of features within the same class even after performing an attack, leading to poor attack performance on new target models. In contrast, the proposed method reduces feature similarity and exhibits superior attack performance compared to conventional methods.
B Extended Experimental Results
To further prove the superiority of the proposed method, we conducted comparative experiments with various conventional methods [36, 17, 1, 53, 32, 54, 11]. To evaluate the transferability of the attack methods, we designed the experiments with non-overlapping encoders for the source model and target model. As shown in Table 1 and  2, the proposed method achieves the best performance among black-box attack methods on the source model (CosPGD [1] is a white-box attack method) and shows superior attack performance on target models compared to existing methods.
Table 1: Attack performance comparison on Pascal VOC 2012 in terms of mIoU. Lower mIoU means better performance and bold numbers denote the best mIoU values for each experimental setup
[…]
Source Models Attack Method Source Model PSPRes101 DV3Res101 FCNVGG16
Clean Images 80.22/80.18 78.39 82.88 59.80
PSPRes50 PGD [36] 7.72 54.73 59.41 45.70
SegPGD [17] 5.41 54.10 58.95 45.43
CosPGD [1] 1.84 56.63 64.37 45.99
DAG [53] 65.82 62.67 66.22 38.91
NI [32] 7.71 33.49 38.52 32.94
DI [54] 6.41 32.00 35.25 37.34
TI [11] 18.28 64.50 69.60 36.80
FSPGD (Ours) 3.39 22.24 16.84 19.75
DV3Res50 PGD [36] 9.74 52.96 56.35 46.39
SegPGD [17] 7.26 52.05 56.50 46.23
CosPGD [1] 1.67 56.82 61.36 45.94
DAG [53] 66.78 62.12 66.84 38.77
NI [32] 9.89 33.86 36.85 34.92
DI [54] 7.35 31.93 32.93 38.30
TI [11] 19.34 64.99 69.80 37.65
FSPGD(Ours) 3.44 21.89 16.57 19.36
Source Models Attack Method Source Model PSPRes50 DV3Res50 FCNVGG16
Clean Images 78.39/82.88 80.22 80.18 59.80